# How segmented is my network?

### v3

@@ Abstract
Network segmentation is a popular security practice for limiting lateral movement, yet practitioners lack a metric to measure how segmented a network actually is. We introduce the first statistically principled metric for network segmentedness based on global edge density, enabling practitioners to quantify what has previously been assessed only qualitatively. Then, we derive a normalized estimator for segmentedness and evaluate its uncertainty using confidence intervals. For a 95% confidence interval with a margin-of-error of $\pm 0.1$, we show that a minimum of $M=97$ sampled node pairs is sufficient. This result is independent of the total number of nodes in the network, provided that node pairs are sampled uniformly at random. We evaluate the estimator through Monte Carlo simulations on Erdős–Rényi, stochastic block models, and real-world enterprise network datasets, demonstrating accurate estimation and well-behaved coverage. Finally, we discuss applications of the estimator, such as baseline tracking, zero trust assessment, and merger integration.
@@ I Introduction
In this paper, we model a network as a graph and study segmentedness as a measurable property of that graph. We show that a simple statistic derived from the network’s global edge density captures meaningful information about network flatness. We provide a principled method to estimate this quantity from observed data and attach confidence intervals to the estimate. Our approach is intentionally lightweight and does not require detailed knowledge of application semantics or traffic intent. Through analysis and simulation, we demonstrate that the estimator is accurate and well-behaved across a range of network models. We argue that this provides security practitioners with a practical, interpretable metric for day-to-day use.
The rest of the paper is organized as follows. Section II reviews prior work on network segmentation and graph-based metrics. Section III introduces our network model and formally defines segmentedness in terms of global edge density. Section IV presents our estimator and derives confidence intervals for segmentedness from observed network data. Section V provides example sample size calculations for practical deployment. Section VI evaluates the estimator using synthetic networks as well as enterprise networking data. Section VII discusses some operational use cases using the segmentedness metric. Section VIII discusses practical considerations for sampling in operational environments. Section IX makes explicit the limitations of our approach. Section X concludes with a summary of our findings.
@@ II-B Network Segmentation Research
In summary, while recent literature has transitioned from qualitative best practices to quantitative evaluation, a significant gap remains. The most advanced current approaches, such as those in [3] and [26], provide measures of risk such as exposure, robustness, and lateral-movement susceptibility. However, these metrics focus on attack outcomes, estimating how likely specific attacks are to succeed in a given network. To our knowledge, no prior work has proposed a scalar metric that quantifies “segmentedness” as an intrinsic topological property—independent of specific threat models or assumptions about how likely individual services are to be compromised. Such a metric would allow practitioners to answer the fundamental question, “How segmented is my network?” by providing a direct measure of network segmentedness (flatness) that can be tracked consistently across different designs and over time.
@@ III Formal Model
We acknowledge that the structural arrangement of permitted connections can influence security outcomes, for example, by creating chokepoints or enabling indirect communication paths. However, as documented in Section II, security practitioners currently lack a simple quantitative metric that summarizes how permissive a network’s communication policy is; our objective prioritizes a single, easily interpretable scalar over metrics that emphasize structure. This objective explicitly rules out structure-oriented measures such as modularity and the Fiedler value, which capture how connections are arranged rather than how much communication is permitted overall. 00footnotemark: 0
00footnotetext: See Appendix A for a comparison of edge density, modularity and the Fiedler value on graphs derived from the Cisco Secure Workload dataset. The Cisco Secure Workload dataset is discussed in Section VI-C.
@@ III-A Network Model
III-B Metric Scope
III-C Flatness and Segmentedness
@@ III-D Properties
  • Interpretability: Segmentedness has a probabilistic interpretation: $S(G)$ represents the probability that two uniformly random nodes cannot communicate directly.
@@ III-E Expected Number of Reachable Neighbors
We can show that for any fixed graph $G$ with edge density $F(G)$, the expected number of directly reachable neighbors of a randomly chosen node is determined solely by the edge density and the number of nodes in the network.
Select a node $v\in V$ uniformly at random, and let $N(v)$ denote the set of nodes that are directly reachable from $v$, i.e., the neighbors of $v$.
For any other node $u\neq v$, define an indicator random variable
$X_{u}=\begin{cases}1&\text{if }\{u,v\}\in E,\\ 0&\text{otherwise}.\end{cases}$ (3)
Then the degree of $v$ can be written as
$|N(v)|=\sum_{u\in V\setminus\{v\}}X_{u}=\deg(v).$ (4)
Since $v$ is selected uniformly at random from $V$, the expected value of $|N(v)|$ equals the average node degree in $G$:
$\mathbb{E}[|N(v)|]=\frac{1}{n}\sum_{w\in V}\deg(w)=\frac{2|E|}{n}.$ (5)
Substituting $|E|=\binom{n}{2}F(G)=\frac{n(n-1)}{2}F(G)$ yields
$\mathbb{E}[|N(v)|]=(n-1)\,F(G).$ (6)
Thus, the expected number of directly reachable neighbors of a randomly selected node is proportional to the edge density and grows linearly with the size of the network.
@@ IV-D Statistical Guarantees and Sample Size
where $z_{\alpha/2}$ is the corresponding quantile of the standard normal distribution. This is known as the Wald or normal approximation confidence interval [1]. 00footnotemark: 0
00footnotetext: We use the Wald confidence interval because it yields a simple, closed-form sample size requirement and enables conservative, network-size-independent guarantees under worst-case variance assumptions.
@@
Because segmentedness is defined as S(G)=1F(G)S(G)=1-F(G), the same confidence interval and error bounds apply directly to $\hat{S}$. Importantly, the required sample size depends only on the desired accuracy and confidence level, and not on the total number of nodes in the network, provided that node pairs are sampled uniformly at random. 00footnotemark: 0
00footnotetext: Compared to edge density, modularity and the Fiedler value (Section II-C) are significantly more challenging to estimate via sampling.
@@ V Example Sample Size Calculations
Rounding up to the nearest integer, this gives a minimum of $M=97$ sampled node pairs. Notably, this result is independent of the total number of nodes in the network, provided that node pairs are sampled uniformly at random.
Table I summarizes the corresponding conservative estimates of $M$ for networks with 10,000 and 100,000 nodes under the same confidence requirements. Tables II and III show the estimates for $M$ for networks with 1,000, 10,000, and 100,000 nodes, when the confidence coefficient and the half-width are varied.
That a 95% CI with a half-width of $\pm 0.1$ can be achieved with just $97$ node pairs is an elegant result—it makes the use of edge density as a measure of segmentedness feasible for networks of arbitrary size. Segmentedness measurements can be repeated frequently (for example, every quarter), even if they have to be run manually or over low-bandwidth links. 00footnotemark: 0
00footnotetext: See Appendix B for handling of the special case when the empirical number of connections is zero.
@@ VI-A Estimation over Random Networks
Figure 1 illustrates the Monte Carlo behavior of the estimator $\hat{F}$ for Erdős–Rényi graphs with n=100,000n=100{,}000 nodes, chosen to mimic the scale of a large organizational network. The true edge probability $p$ is varied over the range $0.1\leq p\leq 0.5$, corresponding to different levels of network connectivity. For each value of $p$, the markers show the mean estimate of $\hat{F}$ across repeated (1,0001{,}000) trials, and the shaded region indicates the spread of estimates across Monte Carlo simulations. The estimates lie close to the diagonal, on either side, across all values of $p$, confirming that the estimator is unbiased in this setting and behaves stably. 00footnotemark: 0
00footnotetext: The diagonal represents perfect estimation: estimate = true value.
@@ VI-B Estimation over Networks with Communities
where the weights $\frac{1}{K}$ and $\frac{K-1}{K}$ represent the proportions of within-block and between-block node pairs, respectively.
@@ VI-C Estimation over Enterprise Network Data
In a well-segmented environment, graphs representing policy-permitted communication will generally be denser than those built from observed traffic over a few days. This is because many permitted pathways, such as disaster recovery routes or rarely used management ports, may remain idle during a short observation period. Indeed, the majority of the graphs in the Cisco dataset exhibit extreme sparsity, with global edge densities $F(G)$ typically falling well below 0.05. 00footnotemark: 0
For this evaluation, we selected the five densest graphs in the dataset to approximate real-world networks and network security policies. These are the $g1$, $g19$, $g21$, $g7$ and $g11$ enterprise networks. As shown in Table IV, the empirical mean edge density $\hat{F}$ calculated over 1,000 Monte Carlo trials matches the true edge density $F(G)$ closely. The true density remains within the calculated 95% confidence intervals for both graphs, empirically validating the estimator’s accuracy and its independence from the total node count $n$.
@@
Having validated the estimator’s accuracy across synthetic and real-world networks, we now turn to practical applications of the segmentedness metric in operational environments.
00footnotetext: In calculating true edge density of these networks, we count any communication between two nodes as evidence of a link between the two nodes.
@@
VII Operational Use Cases
@@
Merger and acquisition integration During mergers or acquisitions, organizations may be required to integrate networks The segmentedness metric provides a quantitative proxy for network security during the integration process. For instance, two networks with pre-merger segmentedness values of $\hat{S_{1}}=0.5$ and $\hat{S_{2}}=0.4$ might temporarily show $\hat{S}_{combined}=0.3$ immediately after connection, reflecting increased connectivity between previously isolated systems. Tracking this metric helps security teams ensure that integration does not inadvertently create an overly flat combined network.
@@
A step-by-step procedure for implementing segmentedness measurements in operational environments is provided in Appendix C.
VIII Practical Considerations for Sampling
The segmentedness metric assumes that node pairs are sampled uniformly at random from the set of networked entities. In operational environments, implementing such sampling introduces practical challenges that may affect how closely real-world measurements adhere to this assumption. While Section IX discusses sampling limitations from a statistical perspective, this section focuses on operational challenges that practitioners may encounter when deploying the measurement methodology.
@@ VIII Practical Considerations for Sampling
IX Limitations
The proposed metric and methodology have several important limitations that should be considered when interpreting results.
@@ IX Limitations
Sampling limitations The statistical guarantees assume uniform random sampling and independent connectivity tests. Non-uniform sampling strategies—such as oversampling from specific subnets or sampling geographically clustered nodes—could introduce correlation between samples and bias estimates. Organizations should ensure sampling is truly random across the entire endpoint population. See Section VIII for a detailed discussion of practical sampling challenges in operational environments.
@@ Appendix A Graph-based Measures’ Comparison
Appendix B Bayesian Inference for High-segmentedness
The use of the Wald confidence interval in (12) poses a mathematical challenge when the number of observed connections is zero. In such cases, the empirical flatness $\hat{F}$ becomes zero, which causes the standard error to collapse and results in a zero-width confidence interval. This ”zero-width” interval is misleading for security practitioners as it suggests absolute certainty that no communication paths exist, when in reality it is merely a result of the empirical sample and the sample size used.
We incorporate a prior belief that most enterprise networks have an edge density of at least 0.01 (one connection per 100 pairs). This 1% baseline reflects the assumption that even highly segmented enterprise networks maintain minimal operational connectivity (e.g., for monitoring, backup, or administrative access). This prior is modeled using a Beta distribution with hyperparameters $\alpha=1$ and $\beta=99$, which centers our initial expectation at 1%.
$\alpha^{\prime}=\alpha+k=1+0=1$ (16)
$\beta^{\prime}=\beta+M=99+97=196$ (17)
Appendix C Deployment Guide
@@
Step 2: Sample node pairs. Generate $M$ random unordered pairs of distinct nodes. For a 95% confidence interval with $\pm 0.1$ half-width, set $M=97$. Use uniform random sampling without replacement from the set of all possible pairs.
@@
Step 5: Compute confidence interval. The 95% Wald confidence interval is $\hat{S}-0.1,\hat{S}+0.1]$. For the special case where $k=0$, use the Bayesian approach described in Appendix B to avoid a zero-width interval.

### v4

@@ Abstract
Network segmentation is a popular security practice for limiting lateral movement, yet practitioners lack a metric to measure how segmented a network actually is. We define segmentedness as the fraction of potential node-pair communications disallowed by policy—equivalently, the complement of graph edge density—and show it to be the first statistically principled scalar metric for this purpose. Then, we derive a normalized estimator for segmentedness and evaluate its uncertainty using confidence intervals. For a 95% confidence interval with a margin-of-error of $\pm 0.1$, we show that a minimum of $M=97$ sampled node pairs is sufficient. This result is independent of the total number of nodes in the network, provided that node pairs are sampled uniformly at random. We evaluate the estimator through Monte Carlo simulations on Erdős–Rényi, stochastic block models, and real-world enterprise network datasets, demonstrating accurate estimation. Finally, we discuss applications of the estimator, such as baseline tracking, zero trust assessment, and merger integration.
@@ I Introduction
In this paper, we model a network as a graph and study segmentedness as a measurable property of that graph. We show that a simple statistic derived from the network’s global edge density captures meaningful information about network flatness. We provide a principled method to estimate this quantity from observed data and attach confidence intervals to the estimate. Our approach is intentionally lightweight and does not require detailed knowledge of application semantics or traffic intent. Through analysis and simulation, we demonstrate that the estimator is accurate and well-behaved (unbiased with correct coverage) across a range of network models. This provides security practitioners with a practical, interpretable metric for day-to-day use.
The rest of the paper is organized as follows. Section II reviews prior work on network segmentation and graph-based metrics. Section III introduces our network model and formally defines segmentedness in terms of global edge density. Section IV presents our estimator and derives confidence intervals for segmentedness from observed network data. Section V provides example sample size calculations for practical deployment. Section VI evaluates the estimator using synthetic networks as well as enterprise networking data. Section VII extends the estimation framework to handle highly segmented networks where the confidence interval degenerates. Section VIII discusses some operational use cases using the segmentedness metric. Section IX discusses practical considerations for sampling in operational environments. Section X makes explicit the limitations of our approach. Section XI concludes with a summary of our findings.
@@ II-B Network Segmentation Research
Because network segmentation spans multiple disciplines and relatively few researchers study it directly, we review the available literature comprehensively rather than selectively.
@@
In summary, while recent literature has transitioned from qualitative best practices to quantitative evaluation, a significant gap remains. The most advanced current approaches, such as those in [3] and [26], provide measures of risk such as exposure, robustness, and lateral-movement susceptibility. However, these metrics focus on attack outcomes, estimating how likely specific attacks are to succeed in a given network. To our knowledge, no prior work has proposed a scalar metric that quantifies “segmentedness” as an intrinsic graph (network) property—independent of specific threat models or assumptions about how likely individual services are to be compromised. Such a metric would allow practitioners to answer the fundamental question, “How segmented is my network?” by providing a direct measure of network segmentedness (flatness) that can be tracked consistently across different designs and over time.
@@ III Formal Model
We acknowledge that the structural arrangement of permitted connections can influence security outcomes, for example, by creating chokepoints or enabling indirect communication paths. However, as documented in Section II, security practitioners currently lack a simple quantitative metric that summarizes how permissive a network’s communication policy is; our objective prioritizes a single, easily interpretable scalar over metrics that emphasize structure. This objective explicitly rules out structure-oriented measures such as modularity and the Fiedler value, which capture how connections are arranged rather than how much communication is permitted overall (see Appendix A for a comparison of edge density, modularity and the Fiedler value).
@@ III-A Network Model
III-B Flatness and Segmentedness
@@ III-D Properties
  • Interpretability: Segmentedness has a probabilistic interpretation: $S(G)$ represents the probability that two uniformly random nodes cannot communicate directly. The probabilistic interpretation holds when a node-pair is selected uniformly at random from the set of all $\binom{n}{2}$ unordered pairs.
@@ III-E Expected Number of Reachable Neighbors
The expected number of directly reachable neighbors of a randomly chosen node depends solely on the edge density and the number of nodes in the network.
Select a node $v\in V$ uniformly at random, and let $N(v)$ denote the set of nodes that are directly reachable from $v$, i.e., the neighbors of $v$. The expected size of $N(v)$ is the average node degree in $G$:
$\mathbb{E}[|N(v)|]=\frac{1}{n}\sum_{w\in V}\deg(w).$ (3)
In any undirected graph, each edge contributes exactly one to the degree of each of its two endpoints, so the sum of all node degrees equals twice the number of edges: $\sum_{w\in V}\deg(w)=2|E|$. Substituting this identity and $|E|=\binom{n}{2}F(G)=\frac{n(n-1)}{2}F(G)$ yields
$\mathbb{E}[|N(v)|]=\frac{2|E|}{n}=(n-1)\,F(G).$ (4)
@@ IV-D Statistical Guarantees and Sample Size
where $z_{\alpha/2}$ is the corresponding quantile of the standard normal distribution. This is known as the Wald or normal approximation confidence interval [1].
@@
Because segmentedness is defined as S(G)=1F(G)S(G)=1-F(G), the same confidence interval and error bounds apply directly to $\hat{S}$. Importantly, the required sample size depends only on the desired accuracy and confidence level, and not on the total number of nodes in the network, provided that node pairs are sampled uniformly at random. 00footnotemark: 0 Compared to edge density, modularity and the Fiedler value (Section II-C) are significantly more challenging to estimate via sampling.
00footnotetext: Independence holds under sampling with replacement (as stated earlier in the paper). When sampling without replacement, a finite-population correction factor of $\sqrt{1-M/\binom{n}{2}}$ applies to the standard error. For all network sizes considered in this paper, this correction is negligible: even for the smallest graph evaluated ($n=86$, $\binom{n}{2}=3{,}655$), the correction factor exceeds $0.98$.
@@ V Example Sample Size Calculations
Rounding up to the nearest integer, this gives a minimum of $M=97$ sampled node pairs.
Table I summarizes the corresponding conservative estimates of $M$ for networks with 1,000, 10,000 and 100,000 nodes under the same confidence requirements. Tables II and III show the estimates for $M$ for networks with 1,000, 10,000, and 100,000 nodes, when the confidence coefficient and the half-width are varied.
That a 95% CI with a half-width of $\pm 0.1$ can be achieved with just $97$ node pairs is an elegant result—it makes the use of edge density as a measure of segmentedness feasible for networks of arbitrary size. Segmentedness measurements can be repeated frequently (for example, every quarter), even if they have to be run manually or over low-bandwidth links. See Section VII for handling of the special case when the empirical number of connections is zero.
@@ VI-A Estimation over Random Networks
Figure 1 illustrates the Monte Carlo behavior of the estimator $\hat{F}$ for Erdős–Rényi graphs with n=100,000n=100{,}000 nodes, chosen to mimic the scale of a large organizational network. The true edge probability $p$ is varied over the range $0.1\leq p\leq 0.5$, corresponding to different levels of network connectivity. For each value of $p$, the markers show the mean estimate of $\hat{F}$ across repeated (1,0001{,}000) trials, and the shaded region indicates the spread of estimates across Monte Carlo simulations. The estimates lie close to the diagonal, where estimate equals true value, across all values of $p$, confirming that the estimator is unbiased in this setting and behaves stably.
@@ VI-B Estimation over Networks with Communities
where the weights $\frac{1}{K}$ and $\frac{K-1}{K}$ represent the proportions of within-block and between-block node pairs, respectively. 00footnotemark: 0
@@
00footnotetext: The exact expression for $K$ equal-sized blocks of size $n/K$ is $F(G)=\frac{n/K-1}{n-1}\,p_{\text{in}}+\frac{n-n/K}{n-1}\,p_{\text{out}}$, which converges to the stated form as $n\to\infty$. For n=100,000n=100{,}000 and $K=5$, the approximation error is negligible.
@@ VI-C Estimation over Enterprise Network Data
In a well-segmented environment, graphs representing policy-permitted communication will generally be denser than those built from observed traffic over a few days. This is because many permitted pathways, such as disaster recovery routes or rarely used management ports, may remain idle during a short observation period. Indeed, the majority of the graphs in the Cisco dataset exhibit extreme sparsity, with global edge densities $F(G)$ typically falling well below 0.05 (in calculating true edge density of these networks, we count any communication between two nodes as evidence of a link between the two nodes).
Although the Cisco dataset reflects observed traffic rather than configured policy, the statistical properties of the estimator—unbiasedness and coverage—depend only on the sampling procedure and the true edge density of the underlying graph, regardless of whether that graph represents policy or observed communication.
For this evaluation, we selected the five densest graphs in the dataset to approximate real-world networks and network security policies. These are the $g1$, $g11$, $g19$, $g21$, and $g7$ enterprise networks. As shown in Table IV, the empirical mean edge density $\hat{F}$ calculated over 1,000 Monte Carlo trials matches the true edge density $F(G)$ closely. The true density remains within the calculated 95% confidence intervals for all graphs, empirically validating the estimator’s accuracy and its independence from the total node count $n$.
@@
VII Bayesian Inference for High-segmentedness
The use of the Wald confidence interval in Section IV-D poses a mathematical challenge when the number of observed connections is zero. In such cases, the empirical flatness $\hat{F}$ becomes zero, which causes the standard error to collapse and results in a zero-width confidence interval. This ”zero-width” interval is misleading for security practitioners as it suggests absolute certainty that no communication paths exist, when in reality it is merely a result of the empirical sample and the sample size used.
We incorporate a prior belief that most enterprise networks have an edge density of at least 0.01 (one connection per 100 pairs). This 1% baseline reflects the assumption that even highly segmented enterprise networks maintain minimal operational connectivity (e.g., for monitoring, backup, or administrative access). This prior is modeled using a Beta distribution with hyperparameters $\alpha=1$ and $\beta=99$, which centers our initial expectation at 1%. 00footnotemark: 0
$\alpha^{\prime}=\alpha+k=1+0=1$ (14)
$\beta^{\prime}=\beta+M=99+97=196$ (15)
00footnotetext: The posterior is sensitive to the choice of prior: for example, $\text{Beta}(1,49)$ shifts the posterior mean to approximately $0.01$ and the $95\%$ upper bound to approximately $0.03$. We select $\text{Beta}(1,99)$ as a conservative baseline reflecting minimal operational connectivity, but practitioners should adjust the prior to reflect their environment.
VIII Operational Use Cases
@@
Merger and acquisition integration During mergers or acquisitions, organizations may be required to integrate networks. The segmentedness metric provides a quantitative proxy for network security during the integration process. For instance, two networks with pre-merger segmentedness values of $\hat{S_{1}}=0.5$ and $\hat{S_{2}}=0.4$ might temporarily show $\hat{S}_{combined}=0.3$ immediately after connection, reflecting increased connectivity between previously isolated systems. Tracking this metric helps security teams ensure that integration does not inadvertently create an overly flat combined network.
@@
In each case, the metric’s independence from network size and its explicit confidence bounds make repeated measurement practical. A step-by-step procedure for implementing segmentedness measurements in operational environments is provided in Appendix B.
IX Practical Considerations for Sampling
The segmentedness metric assumes that node pairs are sampled uniformly at random from the set of networked entities. In operational environments, implementing such sampling introduces practical challenges that may affect how closely real-world measurements adhere to this assumption. While Section X discusses sampling limitations from a statistical perspective, this section focuses on operational challenges that practitioners may encounter when deploying the measurement methodology.
@@ IX Practical Considerations for Sampling
X Limitations
The preceding section discussed operational challenges in deploying the measurement procedure. This section addresses methodological limitations of the metric and estimator themselves.
@@ X Limitations
Sampling limitations The statistical guarantees assume uniform random sampling and independent connectivity tests. Non-uniform sampling strategies—such as oversampling from specific subnets or sampling geographically clustered nodes—could introduce correlation between samples and bias estimates. Organizations should ensure sampling is truly random across the entire endpoint population. See Section IX for a detailed discussion of practical sampling challenges in operational environments.
@@ Appendix A Graph-based Measures’ Comparison
Appendix B Deployment Guide
@@
Step 2: Sample node pairs. Generate $M$ random unordered pairs of distinct nodes. For a 95% confidence interval with $\pm 0.1$ half-width, set $M=97$. Use uniform random sampling with replacement from the set of all possible pairs.
@@
Step 5: Compute confidence interval. The 95% Wald confidence interval is $\hat{S}-0.1,\hat{S}+0.1]$. For the special case where $k=0$, use the Bayesian approach described in Section VII to avoid a zero-width interval.